OpenStudy (zeig_101):

Angle A=190, Angle B=40, Angle C=-96 List the sides of triangle ABC in order from shortest to longest. *EDIT* Full question:m angle A = 8x –2, m angle B = 2x – 8, and m angleC = 94 – 4x. List the sides of triangle ABC in order from shortest to longest.

OpenStudy (anonymous):

so??

OpenStudy (anonymous):

i don't see what you want...review your question up there

OpenStudy (anonymous):

???????????????

OpenStudy (anonymous):

How can an angle be negative? :/

OpenStudy (anonymous):

its not posible

OpenStudy (kainui):

Don't listen to anyone in this, they don't know what they're talking about. Draw out the triangle and I'll help you along.

OpenStudy (kainui):

All of this is completely possible.

OpenStudy (anonymous):

what????? dont know what were talkin about??????????? really sir

OpenStudy (zeig_101):

m angle A = 8x –2, m angle B = 2x – 8, and m angleC = 94 – 4x. List the sides of triangle ABC in order from shortest to longest.

OpenStudy (zeig_101):

I looked at another article on here that had the first half of this listed, and when i worked out the problems for X, the above angles are what I got.

OpenStudy (anonymous):

Okay, now that's better! You've been given that m angle A = 8x –2, m angle B = 2x – 8, and m angleC = 94 – 4x. Sum of the 3 angles of a triangle is 180 thus, equate m<A + m<B + m<C to 180 Do that, and first find out the value of x, and hence the 3 angles. @zeig_101

OpenStudy (zeig_101):

That is what I did, and I got those angles. Could you try to work it out and see if you get something different?

OpenStudy (anonymous):

What are the 3 angles that you got?

OpenStudy (zeig_101):

A=190, B=40, C=-96

OpenStudy (anonymous):

Eh? What do you get as x?

OpenStudy (zeig_101):

24 ...

OpenStudy (anonymous):

Would you mind showing my your calculations?

OpenStudy (anonymous):

me*

OpenStudy (zeig_101):

wait

OpenStudy (zeig_101):

I think I just spotted one error

OpenStudy (anonymous):

Oh!

OpenStudy (zeig_101):

180=8x-2+2x-8+94-4x 6x+84=180 180-84=96 6x=96 96/6=16 ?

OpenStudy (anonymous):

Perfect :D :P

OpenStudy (zeig_101):

ok let me rework something

OpenStudy (zeig_101):

8(16)-2 2(16)-8 94-4(16) 126 24 30 Am I correct?

OpenStudy (anonymous):

Oh yes, you are :D

OpenStudy (zeig_101):

Ok, but now I still need to know which sides are longest(I don't have a protractor to draw the triangle)

OpenStudy (anonymous):

The side opposite to the largest angle has to be the longest side..

OpenStudy (zeig_101):

Sorry I meant the length of all sides, Is there a formula I can use?

OpenStudy (anonymous):

That's what i mean here, <A is the largest, no? this means, the side opposite to <A is the longest .. Similarly, <B being the smallest angle, makes the side opposite to it the shortest. Do you get it?

OpenStudy (zeig_101):

Ok thanks so, BC is largest, AC is smallest?

OpenStudy (anonymous):

Yep!

OpenStudy (zeig_101):

Thanks! Problem solved!

OpenStudy (anonymous):

:D